OpenStudy (anonymous):

Solve using substitution. 6x^2+3y^2=12 y=-x+2

OpenStudy (anonymous):

plug the second equation, where it is solve for y already....into the first equation.

OpenStudy (anonymous):

6x^2+3(-x+2)^2 = 12 -------------->like this

OpenStudy (anonymous):

solve for x first, then after you get the x plug the number you got for x into either equation and that will help you get the y value.

OpenStudy (anonymous):

This is what I've done so far 6x^2+3(-x+2)^2=12 6x^2+3(x^2-4x+4)=12 6x^2+3x^2-12x+12=12 9x^2-12x=0

OpenStudy (anonymous):

x(9x-12)=0 Now try solving

OpenStudy (anonymous):

x=0, 4/3?

OpenStudy (anonymous):

You got it =)

OpenStudy (anonymous):

Now plug those values in to get the values for y

OpenStudy (anonymous):

Okie dokes! Thank you!

OpenStudy (anonymous):

Yup no problem :)